Please help me solve this problem:
3.7 + 7 = 8.1b- 19.4

Answer by oberobic(2304)

3.7 + 7 = 8.1*b -19.4
.
multiply by 10 to eliminate decimals
.
37 +70 = 81b - 194
.
add 194 to both sides
.
37 +70 +194 = 81b
so
81b = 37+70+194
.
81b = 301
.
divide both sides by 81
.
b = 301/81 = approximately 3.716049382716049
.
Done.